14. Step 2012 A random se e d of 35 bushes per rende r e steps the conced ende
Answer #1

sample mean, xbar = 35.8
sample standard deviation, s = 2.79
sample size, n = 9
degrees of freedom, df = n - 1 = 8

Given CI level is 80%, hence α = 1 - 0.8 = 0.2
α/2 = 0.2/2 = 0.1, tc = t(α/2, df) = 1.397


ME = tc * s/sqrt(n)
ME = 1.397 * 2.79/sqrt(9)
ME = 1.299

CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (35.8 - 1.397 * 2.79/sqrt(9) , 35.8 + 1.397 * 2.79/sqrt(9))
CI = (34.5 , 37.1)